# Words as Trigger Points in Social Media Discussions—A Large-Scale Case Study about UK Politics on Reddit

### v2

@@ 1. Space
Discussions on Reddit are structured in different fora, also called subreddits. Aiming to investigate the different effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 3). We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
2. Semantics
We also compare trigger words with a term that shares a similar meaning, but that does not go against the deeper underlying transgressions responsible for trigger responses. To members of the subreddits in Table 3, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^5] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda.
3. Time
In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 3 are triggered by those only in 2022 but not in the respective earlier years.
3.3.2. Statistical analysis
We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 4) and 48 experiments for H2 (Table 5). For each experiment, we prepare the data in the same way. The treatment/placebo is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread. To test H1, we count the number of posts for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group. We then calculate the share of posts before and after the intervention. For example, if there are 8 posts during the 30 minutes before the trigger word and 12 posts during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention. For experiments with H2, we count the number of the respective animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke, 2009):
@@ 3.3.2. Statistical analysis
$\text{interaction}=\text{trigger\_dummy}\times\text{treatment\_dummy}.interaction = trigger_dummy × treatment_dummy$ .
Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al., 2023). It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time. We investigate the plausibility of the assumption in section 4.3.
3.4. Features
We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We simply use post counts for the first hypothesis (H1) related to user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al., 2019) and are trained on social media corpora, specifically X (Twitter). Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al., 2019).
3.4.1. User engagement (Hypothesis 1)
Our first Hypothesis suggests that trigger words lead to higher user engagement. We simply rely on the number of posts to investigate this. That is, we expect to see an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
3.4.2. Animosity (Hypothesis 2)
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, the presence of negative sentiment, anger, and different forms of hate speech.
@@ 3.4.2. Animosity (Hypothesis 2)
Figure 2. Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
@@
For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) which identifies a post as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019). The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender biases (Marjanovic et al., 2022). We expect trigger words to cause more negative sentiments.
@@
We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022) to assign one or more emotions to each tweet. This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018), covering 11 different emotions [^6] Detailed list of emotions can be found in Appendix A.1). For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3. Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
@@
Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados, 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^7] Detailed list of categories can be found in Appendix A.2). The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al., 2021), women’s rights (Farrell et al., 2019), and health issues (Walter et al., 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
4. Results
In this section, we present the results for the two hypotheses tested in this paper.
4.1. User engagement (Hypothesis 1)
@@
Our initial hypothesis is that if trigger words significantly influence online discussions, there should be more messages after mentioning these terms. The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 4).
@@
In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all the terms we selected. It is largest with Rwanda where we observe a 10%-points (Table 4) increase. To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group. In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate. The lowest value is a 3.2%-point increase for the semantic comparison in Brexit. The largest increase is for NHS in the semantic control group, where we observe a 13.4% point increase.
@@
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the semantic and temporal control groups for Rwanda (Appendix B, Figure 9).
@@
Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our initial hypothesis on the effect of trigger words on online discussions.
@@
Table 5. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2. Animosity (Hypothesis 2)
Our second hypothesis suggests that trigger words may increase the frequency of polarising messages and induce animosity in the discussion.
Figure 4. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
Controversiality.
We analyse the average controversiality of each thread and compare the proportion of controversial threads before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the semantic comparison for Rwanda—a causal effect of 6.6% points.
However, significant differences in controversiality do not appear across all settings. Notably, there seems to be no increase in controversial comments when considering Feminism, and in the temporal and semantic settings for Vaccine and NHS, respectively.
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. A similar trend is seen in both the semantic and temporal control settings, as documented in Appendix B, Figure 8).
@@ 4.2. Animosity (Hypothesis 2)
Figure 5. Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
Sentiment & emotion.
@@
We further investigate the effect trigger words have on sentiment and emotion. Based on our hypothesis, trigger words cause users to express more negatively charged emotions. In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
@@
When considering the negatively charged comments of the treatment and control groups for each trigger word, Table 5 indicates a clear trend. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS. The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and semantics experiments.
@@
Let us consider the vaccine trigger, Figure 2, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group shows more messages than the control group, suggesting that the trigger words affects the expression of negative sentiment.
@@
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison). As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure LABEL:fig:emotion_space_vaccine. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
Hate speech.
@@
As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of hate speech or abusive language for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
@@
We illustrate our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in. We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger. Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
4.3. Testing the Parallel Trends Assumption.

### v3

@@ 1. Space
Discussions on Reddit are structured in different fora, also called subreddits. Aiming to investigate the effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 2). We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
2. Contextual
We also compare trigger words with terms that appear in a similar context, but that does not go against the more profound underlying transgressions responsible for trigger responses in similar contexts. To members of the subreddits in Table 2, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^5] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda. Treatment and control groups are from the same subreddits in the same year.
3. Temporal
In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda”, we compare discussions in subreddits likely to spark reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 2 are triggered by those only in 2022 but not in the respective earlier years. Treatment and control groups share the same subreddits but are from different years.
Statistical analysis
We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 3) and 48 experiments for H2 (Table 5). For each experiment, we prepare the data in the same way. The treatment is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread. To test H1, we count the number of comments for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group. We then calculate the share of comments before and after the intervention. For example, if there are 8 comments during the 30 minutes before the trigger word and 12 comments during the 30 minutes after the trigger word, we normalize the data of this discussion thread to 40% before intervention and 60% after intervention. For experiments with H2, we count the number of comments expressing animosity in a thread and then normalize the data as a share of animosity per thread before and after the trigger word.
We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke 2009):
@@ 3. Temporal
[…]
Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al. 2023). It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time. We investigate the plausibility of the assumption in section 4.3.
3.4 Features
We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We use comment counts for the first hypothesis (H1) about user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalize animosity, we build on various language models that follow the RoBERTa architecture (Liu et al. 2019) and are trained in social media corpora, specifically X (Twitter). Even though comments on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al. 2019).
User engagement (Hypothesis 1)
Our first hypothesis suggests that trigger words lead to higher user engagement. We rely on the number of comments to investigate this and expect an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
Animosity (Hypothesis 2)
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalize this with controversial comments, negative sentiment, anger, and different forms of hate speech.
@@ 3.4 Features
Figure 2: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
@@
For sentiment analysis, we use twitter-roberta-base-sentiment-latest (Loureiro et al. 2022) which identifies a comment as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019). As we expect trigger words to cause more negative sentiments, we use sentiment analysis to investigate this, similarly to work using it to gain insights into how controversial topics such as vaccination (Melton et al. 2021), politics (Guimaraes et al. 2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022) are discussed online.
@@
We leverage twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al. 2022) to assign one or more emotions to each tweet. This model is trained on the Semeval 2018 Affect in Tweets task (Mohammad et al. 2018), covering 11 different emotions. [^6] Detailed list of emotions can be found in Appendix A.1) For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
@@
Finally, we use the twitter-roberta-base-hate-multi-class model (Antypas and Camacho-Collados 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups. [^7] Detailed list of categories can be found in Appendix A.2 ) The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al. 2021), women’s rights (Farrell et al. 2019), and health issues (Walter et al. 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist and racism for Rwanda. For NHS, Brexit and Vaccine we consider all hateful comments jointly.
3.5 Validation of Annotation Model
The primary aim of our research is to use simple and computationally efficient models that reliably extract signals relevant to our hypotheses. To confirm the validity and robustness of our findings, we implemented an additional validation procedure to ensure consistency across the classifiers employed in our analysis.
We sampled 100 Reddit comments from our data collection, with 20 comments selected from the subreddits associated with each trigger point (Table 2). The collected comments were independently annotated by three experienced NLP and CSS researchers for sentiment, emotion, and the presence of hate speech. [^8] We include the annotation guidelines in Appendix A.4. Each comment was assigned a label based on the majority vote, while disagreements were resolved through discussion. When considering the difficulty of the annotation task, the annotations present above-average agreement scores for sentiment and hate speech, with Fleiss’ $\kappa$ scores of 0.52 and 0.56, respectively. For the multi-label emotion annotation task, where annotators had to choose from a set of 12 labels,the Krippendorff’s alpha score (Krippendorff 2011) reached 0.24, which is in line with other similar multi-label tasks with a high number of labels (Mohammad et al. 2018; Antypas et al. 2022; Ousidhoum et al. 2019).
We fine-tune the base version of BERT-cased, XLNet, and RoBERTa for sentiment analysis, emotion recognition, and hate speech detection on the same datasets as the models we used for our experiments, as explained in Section 3.4 (Rosenthal, Farra, and Nakov 2017; Mohammad et al. 2018; Antypas and Camacho-Collados 2023) to ensure fairness. We further compare our results to those of five pre-trained models publicly available on HuggingFace. Specifically, two sentiment classifiers trained on Reddit data999 reddit-dBERT: https://huggingface.co/akshataupadhye/finetuning-sentiment-model-reddit-data, reddit-XLNET: https://huggingface.co/minh21/XLNet-Reddit-Sentiment-Analysis , hateBERT (Caselli et al. 2021) and one emotion classifier [^10] BERT-emotion: https://huggingface.co/ayoubkirouane/BERT-Emotions-Classifier. All models are evaluated in the aforementioned manually labelled dataset.
Table 4 presents the average micro F1-scores achieved by the models across each task. We opted for the micro F1-score as it provides a more realistic representation of the results, given the limited sample and heavily imbalanced nature of our dataset. [^11] Detailed dataset distribution is available in Appendix A.3 The Twitter-RoBERTa models achieved the highest scores on all tasks.These scores are matched by XLNet-base for sentiment analysis and XLNet for emotion detection, and the fine-tuned models RoBERTa and XLNet for hate speech detection. Overall, the tested models performed comparably well, except for two off-the-shelf sentiment classifiers that achieved scores below 50%.
Sentiment Emotion Hate
t-RoBERTa 0.76 0.50 0.88
RoBERTa 0.75 0.47 0.88
XLNet 0.76 0.45 0.88
BERT 0.74 0.40 0.85
reddit-dBERT 0.45 - -
reddit-XLNet 0.31 - -
BERT-emotion - 0.45 -
hateBERT - - 0.80
Random 0.30 0.20 0.07
Table 4: Average micro-F1 scores of the models tested for the sentiment analysis, emotion, and hate speech detection tasks. Random indicates a naive baseline that randomly selects labels.
4 Results
4.1 User engagement (Hypothesis 1)
@@
According to our first hypothesis, trigger points cause more comments. The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 3).
@@
In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all terms. It is the largest with Rwanda where we observe a 10%-points (Table 3) increase. To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group. In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate. The lowest value is a 3.2%-point increase for the contextual comparison in Brexit. The largest increase is for NHS in the contextual control group, where we observe a 13.4% point increase.
@@
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 2 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the contextual and temporal control groups for Rwanda (Appendix B, Figure 10).
@@
Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our first hypothesis.
@@
Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2 Animosity (Hypothesis 2)
Our second hypothesis suggests that trigger words may increase the frequency of polarised messages and lead to more animosity in the discussion. [^12] An example of such a discussion is presented in Appendix E.
Figure 4: Average controversiality per thread for the treatment and control (space) groups of Rwanda.
Controversiality. We analyse the average controversiality of each thread and compare the proportion of controversial comments before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the contextual comparison for Rwanda—a causal effect of 6.6% points.
However, significant differences in controversiality do not appear across all settings. Notably, there seems to be no increase in controversial comments for Feminism, and in the temporal and contextual settings for Vaccine and NHS, respectively.
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. The contextual and temporal control settings follow a similar trend, as documented in Appendix B, Figure 9.
@@ 4.2 Animosity (Hypothesis 2)
Figure 5: Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
@@
Sentiment & emotion. We further investigate the effect trigger words have on sentiment and emotion. Based on our hypothesis, trigger words cause users to express more negatively charged emotions. In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
@@
Table 5 indicates a clear trend for the negatively charged comments of the treatment and control groups for each trigger word. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the contextual comparison for NHS. The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and contextual experiments.
@@
Let us consider the vaccine trigger, Figure 3, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group displays a noticeably higher volume of messages around the trigger term, suggesting that triggers amplify the expression of negative sentiment compared to the control condition.
@@
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine contextual comparison). As another example, the overall number of comments expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 12, Appendix B. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
@@
Hate speech. As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of toxicity for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and contextual comparison) and Rwanda (space group).
@@
We illustrate our findings in a visualization, where we consider all messages irrespective of the thread in which they are posted. We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger. Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
4.3 Robustness of Results
We conduct a series of robustness checks designed to test the sensitivity of our results to assumptions and alternative specifications.
Testing the Parallel Trends Assumption.